# Reconstruction of Quantum Fields

### v2

@@ IV Creation–annihilation algebras and transfields
We now turn to the construction of the full creation–annihilation algebra corresponding to a given quadratic quotient of the tensor algebra consistent with transtatistics.
@@ IV.1 Annihilation algebra
$H\otimes H\;=\;R_{\mathrm{gen}}\;\oplus\;R_{\mathrm{gen}}^{\perp},\qquad P_{\mathrm{gen}}+P_{\mathrm{gen}}^{\perp}=\mathrm{id},$
where $R_{\mathrm{gen}}=\mathrm{im}\,P_{\mathrm{gen}}$ is the space of quadratic relations and $R_{\mathrm{gen}}^{\perp}$ is the space of admissible ordered quadratic monomials. Under the dual pairing, this orthogonal decomposition induces a corresponding decomposition
$H^{\ast}\otimes H^{\ast}\;=\;R_{\mathrm{gen}}^{\ast}\;\oplus\;(R_{\mathrm{gen}}^{\perp})^{\ast},$
where $R_{\mathrm{gen}}^{\ast}$ is the annihilator of $R_{\mathrm{gen}}^{\perp}$ and is precisely the image of the adjoint projector
$P_{\mathrm{gen}}^{\ast}\colon H^{\ast}\otimes H^{\ast}\longrightarrow H^{\ast}\otimes H^{\ast},$
@@
$\langle P_{\mathrm{gen}}^{\ast}(\eta),\,v\rangle\;=\;\langle\eta,\,P_{\mathrm{gen}}(v)\rangle,\qquad\eta\in H^{\ast}\otimes H^{\ast},\;v\in H\otimes H.$
@@
$R_{\mathrm{gen}}^{\ast}=\mathrm{im}\,P_{\mathrm{gen}}^{\ast},$
@@
$B\;:=\;T(H^{\ast})\big/\langle R_{\mathrm{gen}}^{\ast}\rangle.$
This construction shows that the creation–annihilation algebras are related by orthogonality: the projector $P_{\mathrm{gen}}$ selects which quadratic combinations of creation operators vanish, while its orthogonal complement $P_{\mathrm{gen}}^{\perp}$ selects the ordered monomials that survive. Under the dual pairing these roles are reversed, so the annihilation relations mirror the creation relations with arrows reversed. In particular, the combinatorial structure governing annihilation is determined by the orthogonality between $R_{\mathrm{gen}}$ and $R_{\mathrm{gen}}^{\perp}$.
@@ IV.2 The Creation–Annihilation Algebras
Theorem 3 (Borowiec–Marcinek, BM2000).
@@ IV.5 Transfield representation of $\mathfrak{gl}(d)$
Using the above relations we can now construct a representation of the general linear algebra in terms of the creation and annihilation operators defining our transtatistical fields (“transfields”). We introduce a family of operators $J_{ij}$ defined using the internal Hermitian form $g$. Let $g_{\alpha\beta}$ be a nondegenerate Hermitian form on the internal space $K$, and let $g^{\beta\alpha}$ denote the entries of its inverse, so that $\sum_{\beta}g_{\alpha\beta}g^{\beta\gamma}=\delta^{\gamma}_{\alpha}$. We define
@@
where $i,j=1,\dots,d$ label the external modes and $\alpha,\beta$ the internal degrees of freedom. We assume that the Hermitian form $g$ is compatible with the exchange maps $R$, $S$, and $C$ in the sense that $C$ is adjoint to $S$ with respect to the inner product on $K\otimes K$, and adjoint to $R$ with respect to the inner product on $K^{*}\otimes K^{*}$, both induced by $g$. Under these adjointness conditions, we may state the following lemma.
@@ V An explicit finite–order example
and let $R_{\mathrm{gen}}:=h^{\perp}\subset K\otimes K$ generate the quadratic relations for a single mode.
@@
With this, we can finally write the relations that generate the algebra. For each mode $i\in\{1,2\}$ the quadratic ideal generated by $R_{\mathrm{gen}}=h^{\perp}\subset K\otimes K$ is given by
@@ VI Conclusions and outlook
On the other hand, although we have defined the quadratic relations that determine the Fock space through the creation operators, the inner product on $\mathcal{H}$ allows us to introduce the annihilation operators in the standard way, namely as the adjoints of the creation operators. In the same way, we use this notion of orthogonality to formalize the creation–annihilation algebras. Moreover, this framework connects with existing work on parastatistics, which share the guiding idea of scrambling internal degrees of freedom that are operationally inaccessible, where we find the classical examples Green53 together with more recent developments WangHazzardProvided. In those cases the relations are introduced heuristically and by postulate, whereas here we obtain a complete characterization derived from basic operational assumptions about the quantum physics of many-particle systems.
A natural generalisation of this work is its application to quantum field theory. Since we have introduced new commutation relations, one can move from discrete degrees of freedom to measure spaces such as those modelling spacetime. This would allow the description of new types of fields compatible with transtatistics, while remaining consistent with the operational arguments presented here regarding accessible information and operational indistinguishability, but now in the continuum. Of particular interest are the consequences of imposing a causal structure on the accessible degrees of freedom, as this additional symmetry could further constrain, or even fully collapse, the Yang–Baxter symmetry of the projectors to the standard bosonic and fermionic cases. This can be viewed as an extension of the spin–statistics theorem within the present framework Pauli1940; StreaterWightman1964; Greenberg1991.
@@ Acknowledgements
Appendix A Lemma 1
Proof.
@@ Appendix A Lemma 1
Appendix B Lemma 2
Proof.
@@ Appendix B Lemma 2
Appendix C Lemma 3
Proof.
@@ Appendix C Lemma 3
Appendix D Theorem 1
This is based on PolishchukPositselski2005.
1. PBW Bases
@@ 2. PBW Theorem
Appendix E Theorem 2
We recall the partition theorem MedinaSanchezDakic2024
@@ Appendix E Theorem 2
By construction: the ordered basis conditions implies the factorization property of the $U(1)$ character and additionally, the $U(d)$ invariance completes the conditions for the application of the partition theorem. The result follows. ∎
Appendix F Lemma 4

### v3

@@ IV Creation–annihilation algebras and transfields
The preceding analysis has determined the algebraic skeleton of the many-particle space: a quadratic quotient of the tensor algebra, constrained by unitarity and the PBW condition, whose single-mode Hilbert-Poincare series characterises the statistics completely. What this construction provides, however, is only the space of states together with the action of the creation operators. A full quantum mechanical framework requires more: one needs annihilation operators, their commutation relations with the creators, and a representation of the observable algebra on the resulting Fock space.
The passage from the creation algebra to the full creation-annihilation algebra is not automatic. It requires specifying how an annihilation operator moves past a creation operator, and this exchange rule must be compatible with the quadratic relations already imposed on each sector separately. The data that controls this compatibility is a linear map between the mixed tensor products of the one-particle space and its dual, which we call the cross map. Its admissible forms are constrained by Yang-Baxter type identities inherited from the PBW condition, and its non-homogeneous part is fixed by the vacuum two-point function. The present section develops this construction in full and concludes by showing that the resulting algebra carries a natural representation of the general linear group, providing the transfield analogue of the standard Fock representation.
@@ IV.1 Annihilation algebra
Throughout Section IV we fix the quadratic data determined in Section III. Namely, let
$R:=\operatorname{im}P_{g}\subset H\otimes H,\qquad R^{\perp}:=\ker P_{g},$
so that the creation algebra is
$A:=T(H)/\langle R\rangle.$
We define the dual quadratic subspace
$S:=R^{*}=\operatorname{im}P_{g}^{*}\subset H^{*}\otimes H^{*},$
and the annihilation algebra
$B:=T(H^{*})/\langle S\rangle.$
Finally, we fix a linear map
$C:H^{*}\otimes H\to H\otimes H^{*}$
(the cross map) which will encode the exchange relations between annihilation and creation operators and is required to satisfy the compatibility conditions of Theorem 3.
@@
$H\otimes H\;=\;R_{\mathrm{g}}\;\oplus\;R_{\mathrm{g}}^{\perp},\qquad P_{\mathrm{g}}+P_{\mathrm{g}}^{\perp}=\mathrm{id},$
where $R_{\mathrm{g}}=\mathrm{im}\,P_{\mathrm{g}}$ is the space of quadratic relations and $R_{\mathrm{g}}^{\perp}$ is the space of admissible ordered quadratic monomials. Under the dual pairing, this orthogonal decomposition induces a corresponding decomposition
$H^{\ast}\otimes H^{\ast}\;=\;R_{\mathrm{g}}^{\ast}\;\oplus\;(R_{\mathrm{g}}^{\perp})^{\ast},$
where $R_{\mathrm{g}}^{\ast}$ is the annihilator of $R_{\mathrm{g}}^{\perp}$ and is precisely the image of the adjoint projector
$P_{\mathrm{g}}^{\ast}\colon H^{\ast}\otimes H^{\ast}\longrightarrow H^{\ast}\otimes H^{\ast},$
@@
$\langle P_{\mathrm{g}}^{\ast}(\eta),\,v\rangle\;=\;\langle\eta,\,P_{\mathrm{g}}(v)\rangle,\qquad\eta\in H^{\ast}\otimes H^{\ast},\;v\in H\otimes H.$
@@
$R_{\mathrm{g}}^{\ast}=\mathrm{im}\,P_{\mathrm{g}}^{\ast},$
@@
$B\;:=\;T(H^{\ast})\big/\langle R_{\mathrm{g}}^{\ast}\rangle.$
This construction shows that the creation–annihilation algebras are related by orthogonality: the projector $P_{\mathrm{g}}$ selects which quadratic combinations of creation operators vanish, while its orthogonal complement $P_{\mathrm{g}}^{\perp}$ selects the ordered monomials that survive. Under the dual pairing these roles are reversed, so the annihilation relations mirror the creation relations with arrows reversed. In particular, the combinatorial structure governing annihilation is determined by the orthogonality between $R_{\mathrm{g}}$ and $R_{\mathrm{g}}^{\perp}$.
@@ IV.2 The Creation–Annihilation Algebras
Theorem 3 (Borowiec–Marcinek, Borowiec and Marcinek (2000)).
@@
Here $R\subset H\otimes H$ and $S\subset H^{*}\otimes H^{*}$ denote the quadratic relation subspaces defining the creation and annihilation algebras, while $C:H^{*}\otimes H\to H\otimes H^{*}$ is the cross map introduced above.
@@ IV.4 The $(A,B)$ bracket
A useful special case is obtained when the tensors $A$ and $B$ arise from a single operator
$R:K\otimes K\to K\otimes K$
by
$A=\frac{1-R}{2},\qquad B=\frac{1+R}{2}.$
Then $A+B=\mathrm{id}$ and AB=RA-B=-R, so the $(A,B)$-bracket takes the form
$P_{i\alpha},Q_{j\beta}]_{A,B}=P_{i\alpha}Q_{j\beta}-\sum_{\eta,\lambda}R^{\eta\lambda}_{\alpha\beta}\,Q_{j\lambda}P_{i\eta}.$
Accordingly, the defining relations become
$X^{\dagger}_{i\alpha}X^{\dagger}_{j\beta}=\sum_{\eta,\lambda}R^{\eta\lambda}_{\alpha\beta}\,X^{\dagger}_{j\lambda}X^{\dagger}_{i\eta},\qquad X_{i\alpha}X_{j\beta}=\sum_{\eta,\lambda}R^{\eta\lambda}_{\alpha\beta}\,X_{j\lambda}X_{i\eta},$
and
$X_{i\alpha}X^{\dagger}_{j\beta}=\sum_{\eta,\lambda}R^{\eta\lambda}_{\alpha\beta}\,X^{\dagger}_{j\lambda}X_{i\eta}+g_{\alpha\beta}\delta_{ij}\mathbf{1}.$
After a relabelling of internal indices, this is precisely the form of the R-paraparticle exchange relations Wang and Hazzard (2025). That framework is therefore recovered as a special case of the present bracket formalism, namely when the quadratic data are encoded by a single involutive Yang–Baxter operator $R$.
@@ IV.5 Transfield representation of $\mathfrak{gl}(d)$
Using the above relations we can now construct a representation of the general linear algebra in terms of the creation and annihilation operators defining our transtatistical fields (namely Transfields). We introduce a family of operators $J_{ij}$ defined using the internal Hermitian form $g$. Let $g_{\alpha\beta}$ be a nondegenerate Hermitian form on the internal space $K$, and let $g^{\beta\alpha}$ denote the entries of its inverse, so that $\sum_{\beta}g_{\alpha\beta}g^{\beta\gamma}=\delta^{\gamma}_{\alpha}$. We define
@@
where $i,j=1,\dots,d$ label the external modes and $\alpha,\beta$ the internal degrees of freedom. We say that the Hermitian form $g$ on $K$ is compatible with the quadratic data $(R,S,C)$ if the projector $P_{g}$ is self-adjoint with respect to the induced inner product on $H\otimes H$, and the cross map $C$ is adjoint to the quadratic data under the pairing induced by $g$. Under these adjointness conditions, we may state the following lemma.
@@ V An explicit finite–order example
and let $R_{\mathrm{g}}:=h^{\perp}\subset K\otimes K$ generate the quadratic relations for a single mode.
@@
With this, we can finally write the relations that generate the algebra. For each mode $i\in\{1,2\}$ the quadratic ideal generated by $R_{\mathrm{g}}=h^{\perp}\subset K\otimes K$ is given by
@@ VI Conclusions and outlook
On the other hand, although we have defined the quadratic relations that determine the Fock space through the creation operators, the inner product on $\mathcal{H}$ allows us to introduce the annihilation operators in the standard way, namely as the adjoints of the creation operators. In the same way, we use this notion of orthogonality to formalize the creation–annihilation algebras. Moreover, this framework connects with existing work on parastatistics, which share the guiding idea of scrambling internal degrees of freedom that are operationally inaccessible, where we find the classical examples Green (1953) together with more recent developments Wang and Hazzard (2025). In those cases the relations are introduced heuristically and by postulate, whereas here we obtain a complete characterization derived from basic operational assumptions about the quantum physics of many-particle systems.
From a broader mathematical perspective, the present construction contributes to the problem of identifying Koszul algebras whose Hilbert functions are totally positive (Pólya frequency) sequences, as formulated in Reiner and Welker (2005). Since PF sequences coincide with totally positive sequences, this question can be viewed as a structural problem at the interface of commutative algebra, combinatorics, and representation theory. Recent work Sam and Vandebogert (2024) further highlights deep connections between total positivity and algebraic constructions such as Schur-type modules and free resolutions.
The class of PBW $U(d)$-equivariant algebras introduced here provides a concrete and physically motivated family for which total positivity holds and is completely controlled by quadratic data. This suggests that the operational principles underlying indistinguishability naturally single out a rigid subclass of Koszul-type algebras in which total positivity is not merely a property, but a consequence of the underlying algebraic and symmetry structure.
A natural generalisation of this work is its application to quantum field theory. Since we have introduced new commutation relations, one can move from discrete degrees of freedom to measure spaces such as those modelling spacetime. This would allow the description of new types of fields compatible with transtatistics, while remaining consistent with the operational arguments presented here regarding accessible information and operational indistinguishability, but now in the continuum. Of particular interest are the consequences of imposing a causal structure on the accessible degrees of freedom, as this additional symmetry could further constrain, or even fully collapse, the Yang–Baxter symmetry of the projectors to the standard bosonic and fermionic cases. This can be viewed as an extension of the spin–statistics theorem within the present framework Pauli (1940); Streater and Wightman (1964); Greenberg (1991).
@@ Acknowledgements
Conflict of Interest
The author declares that there is no conflict of interest.
Data Availability
No data were generated or analyzed in this study.
Appendix A Algebraic preliminaries
A.1 Ordered basis implies quadratic generation
Lemma 1.
@@ A.1 Ordered basis implies quadratic generation
A.2 Factorisation as a tensor power
Lemma 2.
@@ A.2 Factorisation as a tensor power
A.3 $U(d)$-invariant quadratic subspaces
Lemma 3.
@@ A.3 $U(d)$-invariant quadratic subspaces
Appendix B PBW property and Yang–Baxter constraints
This is based on Polishchuk and Positselski (2005).
PBW Bases
@@ Appendix B PBW property and Yang–Baxter constraints
Appendix C Single-mode structure and proof of the main theorem
C.1 Construction of the spaces $W_{n}$
Let $K$ be the internal space and let
$P_{\mathrm{g}}:K\otimes K\longrightarrow K\otimes K$
be the quadratic projector determining the relations. For each $n\geq 0$ define
$V_{n}:=K^{\otimes n}.$
For $1\leq k\leq n-1$ we introduce the operators
$P^{(n)}_{k}:=\mathrm{id}^{\otimes(k-1)}\otimes P_{\mathrm{g}}\otimes\mathrm{id}^{\otimes(n-k-1)}:V_{n}\longrightarrow V_{n},$
acting nontrivially only on the $k$-th and $(k+1)$-th tensor factors.
The space of admissible degree-$n$ monomials is then defined as
$W_{n}:=\bigcap_{k=1}^{n-1}\ker P^{(n)}_{k}\subseteq K^{\otimes n}.$
Equivalently, $W_{n}$ consists of those tensors in $K^{\otimes n}$ that survive all quadratic reductions, i.e. those configurations for which no adjacent pair lies in the generating subspace $\mathrm{im}\,P_{\mathrm{g}}$.
By construction, the family $\{W_{n}\}_{n\geq 0}$ defines a graded vector space
$W:=\bigoplus_{n\geq 0}W_{n},$
and the single-mode Hilbert–Poincaré series is given by
$G(t)=\sum_{n\geq 0}(\dim W_{n})\,t^{n}.$
C.2 Factorisation of the Hilbert–Poincaré series
Let $H\cong\mathbb{C}^{d}\otimes K$ and let
$F=T(H)/\langle R_{\mathrm{g}}\rangle$
be the quadratic quotient admitting a PBW basis. By Lemma 2 there exists a graded vector space
$W=\bigoplus_{n\geq 0}W_{n}$
such that
$F\cong W^{\otimes d}$
as graded vector spaces.
For each $n\geq 0$, the homogeneous component $F_{n}$ decomposes as
$F_{n}\cong\bigoplus_{\begin{subarray}{c}n_{1}+\cdots+n_{d}=n\\ n_{i}\geq 0\end{subarray}}W_{n_{1}}\otimes\cdots\otimes W_{n_{d}}.$
Taking dimensions, we obtain
$\dim F_{n}=\sum_{\begin{subarray}{c}n_{1}+\cdots+n_{d}=n\\ n_{i}\geq 0\end{subarray}}\prod_{i=1}^{d}\dim W_{n_{i}}.$
Define the Hilbert–Poincaré series
$H_{F}(t):=\sum_{n\geq 0}(\dim F_{n})\,t^{n},\qquad G(t):=\sum_{n\geq 0}(\dim W_{n})\,t^{n}.$
Then
$H_{F}(t)=\sum_{n\geq 0}\left(\sum_{n_{1}+\cdots+n_{d}=n}\prod_{i=1}^{d}\dim W_{n_{i}}\right)t^{n}=\left(\sum_{m\geq 0}(\dim W_{m})\,t^{m}\right)^{d}=G(t)^{d}.$
This proves that the Hilbert–Poincaré series factorises as
$H_{F}(t)=G(t)^{d},$
where $G(t)$ is the single-mode series.
C.3 Partition Theorem
We recall the partition theorem Sánchez and Dakić (2024)
@@ C.4 $1\rightarrow 2:$
By construction: the ordered basis conditions implies the factorization property of the $U(1)$ character and additionally, the $U(d)$ invariance completes the conditions for the application of the partition theorem. The result follows.
C.5 $2\rightarrow 1:$
Assume
$G(t)=\frac{Q_{-}(t)}{Q_{+}(t)},\qquad Q_{\pm}(t)\in\mathbb{Z}[t],\quad Q_{+}(0)=1,$
and that all roots of $Q_{+}$ are real and strictly positive while all roots of $Q_{-}$ are real and strictly negative. We will construct a homogeneous quadratic algebra $A$ with a PBW basis, equipped with a unitary invariance, whose single-variable Hilbert–Poincaré series equals $G(t)$.
C.5.1 Step 1. Reduction to the transbosonic case.
Replacing $G(t)$ by $G(-t)^{-1}$ corresponds (at the level of Hilbert series of quadratic algebras) to passing to the Koszul dual: if $A$ is a quadratic algebra then the Hilbert series of its quadratic dual $A^{!}$ is obtained by the standard inversion/grade sign change relation for Koszul algebras. Hence it suffices to construct $A$ in the case
$G(t)=\frac{1}{Q_{+}(t)}=\prod_{i=1}^{r}(1-\alpha_{i}t)^{-1},\qquad\alpha_{i}>0,$
with $\alpha_{i}$ nonzero real numbers (not a priori assumed integral). Once a PBW quadratic realization is obtained in this transbosonic form, the general case follows by duality and the evident functoriality of the construction with respect to quadratic duals.
C.5.2 Step 2. Schur-positive specialization via Edrei–Thoma.
Let $\Lambda$ denote the ring of symmetric functions over $\mathbb{Z}$ and let $h_{n}$ be the complete symmetric functions. Define a sequence
$a_{n}:=[t^{n}]\,G(t)\in\mathbb{Z}_{\geq 0},\qquad n\geq 0,$
and consider the linear functional $\phi:\Lambda\to\mathbb{Z}$ determined by $\phi(h_{n})=a_{n}$. The desired conclusion is that $\phi$ is a Schur-positive specialization (i.e. $\phi(s_{\lambda})\in\mathbb{Z}_{\geq 0}$ for every partition $\lambda$).
The correct necessary and sufficient condition for $\phi$ to be Schur-positive is that the sequence $(a_{n})_{n\geq 0}$ be a nonnegative totally positive sequence in the sense of Edrei and Thoma. The Edrei–Thoma classification (together with finite-type rationality results of Aissen–Schoenberg–Whitney) asserts that a nonnegative integer sequence $(a_{n})$ which is totally positive and whose ordinary generating function is a rational function of finite type admits a canonical factorization of the form
$G(t)=\exp(\gamma t)\,\frac{\prod_{j}(1+\beta_{j}t)}{\prod_{i}(1-\alpha_{i}t)},$
with parameters $\gamma,\alpha_{i},\beta_{j}\geq 0$ (counted with multiplicity). Conversely, such a product produces a totally positive sequence Edrei (1953).
In our hypothesis $G(t)=Q_{-}(t)/Q_{+}(t)$ is a rational function with all poles strictly positive and zeros strictly negative. The integrality $a_{n}\in\mathbb{Z}_{\geq 0}$ together with the finite rational form forces the exponential factor to vanish ($\gamma=0$) and forces the pole/zero multiplicities to be integers; thus $G(t)$ can be rewritten as a finite product of factors $(1-\alpha_{i}t)^{-1}$ (with positive real $\alpha_{i}$ repeated according to integer multiplicity) times finite factors $(1+\beta_{j}t)$ corresponding to negative real zeros (the latter will be absorbed into $Q_{-}$). Hence the Edrei–Thoma theory guarantees that $\phi$ is Schur-positive and, moreover, that the values $\phi(s_{\lambda})$ are nonnegative integers (the latter follows from interpreting the specialization as evaluation on a finite virtual alphabet with integer multiplicities in the finite rational case).
C.5.3 Step 3. Construction of an exact monoidal functor $F$.
Let $\mathrm{Rep}_{\mathrm{poly}}(GL)$ be the semisimple tensor category of polynomial representations of $GL$. Its simple objects are indexed by partitions and are the Schur functors $S_{\lambda}$. Define vector spaces $M_{\lambda}$ with
$\dim M_{\lambda}:=\phi(s_{\lambda})\in\mathbb{Z}_{\geq 0},$
and define an additive functor
$F:\mathrm{Rep}_{\mathrm{poly}}(GL)\longrightarrow\mathrm{Vect}$
by $F(S_{\lambda}):=M_{\lambda}$ and extending additively. The compatibility of $\phi$ with the Littlewood–Richardson coefficients,
$s_{\lambda}s_{\mu}=\sum_{\nu}c_{\lambda\mu}^{\nu}s_{\nu}\quad\Longrightarrow\quad\phi(s_{\lambda})\phi(s_{\mu})=\sum_{\nu}c_{\lambda\mu}^{\nu}\phi(s_{\nu}),$
provides the numerical multiplicity data necessary to equip the family $(M_{\lambda})$ with bilinear maps
$M_{\lambda}\otimes M_{\mu}\longrightarrow\bigoplus_{\nu}M_{\nu}^{\oplus c_{\lambda\mu}^{\nu}}$
which can be chosen (by picking explicit models and coherence isomorphisms) so that $F$ becomes a (weak) monoidal functor. Because $\mathrm{Rep}_{\mathrm{poly}}(GL)$ is semisimple, any additive functor defined on simples and extended linearly is exact Fulton and Harris (2004). Therefore $F$ is exact and monoidal, and it respects the grading by polynomial degree.
C.5.4 Step 4. Apply $F$ to the symmetric algebra.
For finite-dimensional vector spaces $U$ and $V$ we have the Cauchy decomposition
$\mathrm{Sym}(U\otimes V)\cong\bigoplus_{\lambda}S_{\lambda}(U)\otimes S_{\lambda}(V).$
Applying $F$ to the $V$-factor yields a graded algebra
$A:=F\big(\mathrm{Sym}(U\otimes V)\big)\cong\bigoplus_{\lambda}S_{\lambda}(U)\otimes M_{\lambda}.$
Because $F$ preserves degrees and dimensions by construction, the degree-$n$ component of $A$ satisfies $\dim A_{n}=a_{n}$, and therefore the Hilbert–Poincaré series of $A$ equals $G(t)$.
C.5.5 Step 5. Quadratic presentation and preservation of Koszulity / PBW.
The algebra $\mathrm{Sym}(U\otimes V)$ is quadratic and Koszul; write it as
$\mathrm{Sym}(U\otimes V)\cong T(H)/\langle R\rangle,$
with $H=(U\otimes V)_{1}$ and $R\subset H\otimes H$ the quadratic relations. Since $F$ is exact and monoidal it sends the degree-1 generating object $H$ to $F(H)$ and the subobject of quadratic relations $R$ to a subspace $F(R)\subset F(H)\otimes F(H)$, yielding a quadratic presentation
$A\cong T\big(F(H)\big)/\langle F(R)\rangle.$
To conclude that $A$ is Koszul and admits a PBW basis we invoke the theory of Koszul algebras in monoidal abelian categories and preservation results for Koszulity under exact monoidal functors: exact monoidal functors that preserve the quadratic presentation carry Koszul complexes to Koszul complexes and preserve the vanishing/ext properties that characterize Koszul algebras. Thus Koszulness and the PBW property descend from $\mathrm{Sym}(U\otimes V)$ to $A$ under our hypotheses; see Polishchuk and Positselski (2005) for the systematic treatment of Koszul algebras, quadratic duality, and preservation under suitable functors.
C.5.6 Step 6. Presentation and unitary invariance.
By standard theory of quadratic algebras there exist $H^{\prime}$ and $R^{\prime}\subset H^{\prime}\otimes H^{\prime}$ with
$A\cong T(H^{\prime})/\langle R^{\prime}\rangle.$
The construction is functorial in the $U$-factor, since $F$ acts only on the $V$-factor; hence the natural $GL(U)$-action on $\mathrm{Sym}(U\otimes V)$ is transported to $A$. Restricting this action to the unitary subgroup $U(d)$ (for $d=\dim U$) yields a $U(d)$-invariant quadratic algebra. Consequently $A$ is a homogeneous quadratic, $U(d)$-invariant algebra with a PBW basis and single-variable Hilbert–Poincaré series $G(t)$, completing the proof. ∎
Appendix D Creation–annihilation structure
D.1 Dual quadratic algebra
Let $H$ be a finite-dimensional complex vector space and
$A:=T(H)/\langle R_{\mathrm{g}}\rangle$
the quadratic algebra defined by the generating subspace
$R_{\mathrm{g}}\subset H\otimes H,\qquad H\otimes H=R_{\mathrm{g}}\oplus R_{\mathrm{g}}^{\perp}.$
Let $H^{*}$ denote the dual space, equipped with the natural pairing
$\langle\cdot,\cdot\rangle:(H^{*}\otimes H^{*})\times(H\otimes H)\to\mathbb{C}.$
We define the dual quadratic subspace as the annihilator of $R_{\mathrm{g}}^{\perp}$:
$R_{\mathrm{g}}^{*}:=\left\{\eta\in H^{*}\otimes H^{*}\;\middle|\;\langle\eta,v\rangle=0\ \forall\,v\in R_{\mathrm{g}}^{\perp}\right\}.$
Equivalently,
$R_{\mathrm{g}}^{*}=(R_{\mathrm{g}}^{\perp})^{\circ}\subset H^{*}\otimes H^{*}.$
If $P_{\mathrm{g}}:H\otimes H\to R_{\mathrm{g}}$ denotes the projector onto the generating subspace, we define its adjoint
$P_{\mathrm{g}}^{*}:H^{*}\otimes H^{*}\to H^{*}\otimes H^{*}$
by
$\langle P_{\mathrm{g}}^{*}(\eta),v\rangle=\langle\eta,P_{\mathrm{g}}(v)\rangle,\qquad\eta\in H^{*}\otimes H^{*},\ v\in H\otimes H.$
Then
$R_{\mathrm{g}}^{*}=\mathrm{im}\,P_{\mathrm{g}}^{*}.$
The dual quadratic algebra (annihilation algebra) is defined as
$B:=T(H^{*})/\langle R_{\mathrm{g}}^{*}\rangle.$
By construction, the quadratic relations of $B$ are orthogonal to the admissible quadratic monomials of $A$, and the pair $(A,B)$ is related by duality at the level of quadratic data.
D.2 Descent of the cross map
Let
$A=T(H)/\langle R\rangle,\qquad B=T(H^{*})/\langle S\rangle$
be quadratic algebras, where $R\subset H\otimes H$ and $S\subset H^{*}\otimes H^{*}$ are the generating subspaces.
Let
$C:H^{*}\otimes H\longrightarrow H\otimes H^{*}$
be a linear map (the cross map), and let
$\tilde{\tau}:T(H^{*})\otimes T(H)\longrightarrow T(H)\otimes T(H^{*})$
be its multiplicative extension, defined recursively by requiring
$\tilde{\tau}(\eta\otimes x)=C(\eta\otimes x),\qquad\eta\in H^{*},\ x\in H,$
and compatibility with concatenation.
In order for $\tilde{\tau}$ to descend to a well-defined map
$\tau:B\otimes A\longrightarrow A\otimes B,$
it is necessary and sufficient that the defining ideals be stable under $\tilde{\tau}$. Concretely, this means:
$\tilde{\tau}(H^{*}\otimes\langle R\rangle)\subseteq\langle R\rangle\otimes T(H^{*}),$
$\tilde{\tau}(\langle S\rangle\otimes H)\subseteq T(H)\otimes\langle S\rangle.$
Equivalently, since the ideals are generated in degree two, it suffices to impose the conditions at the quadratic level:
$\tilde{\tau}(H^{*}\otimes R)\subseteq R\otimes H^{*},$
$\tilde{\tau}(S\otimes H)\subseteq H\otimes S.$
Under these conditions, the map $\tilde{\tau}$ passes to the quotient and defines a bilinear map
$\tau:B\otimes A\to A\otimes B,$
which encodes the exchange relations between annihilation and creation generators.
D.3 Mixed relations and bracket form
Let $A$ and $B$ be the creation and annihilation algebras constructed above, and let
$\tau=\tilde{\tau}+G:H^{*}\otimes H\longrightarrow H\otimes H^{*}\oplus\mathbb{C}$
be the full cross map, where $G:H^{*}\otimes H\to\mathbb{C}$ is the vacuum two-point function.
Fix a basis $\{X_{i\alpha}\}$ of $H$ and dual basis $\{X^{*}_{i\alpha}\}$ of $H^{*}$. The mixed relations are determined by the exchange rule
$X_{i\alpha}X^{\dagger}_{j\beta}=\sum_{\gamma,\delta}C^{\gamma\delta}_{\alpha\beta}\,X^{\dagger}_{j\delta}X_{i\gamma}+g_{\alpha\beta}\,\delta_{ij}\,\mathbf{1},$
where $C$ encodes the homogeneous part of $\tau$ and $g_{\alpha\beta}$ is the Hermitian form on the internal space.
It is convenient to rewrite all quadratic relations in a unified way using a bracket adapted to the internal projectors. Let $A^{\eta\lambda}_{\alpha\beta}$ and $B^{\eta\lambda}_{\alpha\beta}$ be the tensors determined by the symmetric and antisymmetric components of the quadratic data. For any generators $P_{i\alpha},Q_{j\beta}$ we define
$P_{i\alpha},Q_{j\beta}]_{A,B}:=\sum_{\eta,\lambda}A^{\eta\lambda}_{\alpha\beta}\,\{P_{i\eta},Q_{j\lambda}\}+\sum_{\eta,\lambda}B^{\eta\lambda}_{\alpha\beta}\,[P_{i\eta},Q_{j\lambda}],$
where $\{\cdot,\cdot\}$ and $\cdot$$\cdot$] denote the anticommutator and commutator, respectively.
In terms of this bracket, the defining relations take the uniform form
$X^{\dagger}_{i\alpha},X^{\dagger}_{j\beta}]_{A,B}=0,$
$X_{i\alpha},X_{j\beta}]_{A,B}=0,$
$X_{i\alpha},X^{\dagger}_{j\beta}]_{A,B}=g_{\alpha\beta}\,\delta_{ij}\,\mathbf{1}.$
These relations generalise the canonical commutation and anticommutation relations and encode the full quadratic structure of the creation–annihilation algebra in a form adapted to the internal projectors.
D.4 Representation of $\mathfrak{gl}(d)$